# ppt: A Process-based Preference Learning Framework for Self Improving Table Question Answering Models

### v1

@@ Abstract
Improving large language models (LLMs) with self-generated data has demonstrated success in tasks such as mathematical reasoning and code generation. Yet, no exploration has been made on table question answering (TQA), where a system answers questions based on tabular data. Addressing this gap is crucial for TQA, as effective self-improvement can boost performance without requiring costly or manually annotated data. In this work, we propose ppt, a Process-based Preference learning framework for TQA. It decomposes reasoning chains into discrete states, assigns scores to each state, and samples contrastive steps for preference learning. Experimental results show that ppt effectively improves TQA models by up to 5% on in-domain datasets and 2.4% on out-of-domain datasets, with only 8,000 preference pairs. Furthermore, the resulting models achieve competitive results compared to more complex and larger state-of-the-art TQA systems, while being five times more efficient during inference.
@@ 1 Introduction
Table Question Answering (TQA) aims to answer questions based on tabular data. Current approaches developing TQA models Wu and Feng (2024); Zhang et al. (2025b) fine-tune pre-trained large language models (LLMs) on labeled answers augmented with Chain-of-Thought (CoT) reasoning Wei et al. (2023).
Figure 1: Given a TQA problem, a TQA model generates reasoning states, composed of cumulative reasoning steps. A value function is used to score each state. We roll out parent states to obtain pair-wise steps and select those with value differences larger than a specified threshold. After data collection, we perform step-wise preference learning for improving the TQA model.
Recent studies suggest that LLM performance can be improved by fine-tuning with self-generated data on many tasks, such as mathematical reasoning Singh et al. (2024) and code generation Zelikman et al. (2024). Yet, no exploration of self-improvement has been conducted in TQA. Addressing this gap is crucial, as an effective improvement method further enhances an existing model without costly or manually annotated data.
To self-improve TQA models, two challenges arise: (1) devising an effective learning strategy and (2) automatically obtaining high-quality data for learning. Drawing inspiration from contrastive process-based learning Lai et al. (2024); Zhang et al. (2025a), where pairs of intermediate reasoning steps provide contrastive signals for model learning, we propose ppt, a novel Process-based Preference learning framework for TQA.
As shown in Figure 1, our framework contains three stages: state generation, state value estimation and pair-wise step sampling. During the first stage, we sample multiple reasoning chains from a TQA model and formulate them into reasoning states, composed of cumulative reasoning steps, During state value estimation, a value function gives continuous scores based on the potential of a state for deriving a final correct answer. During pair-wise steps sampling, we roll out parent states to obtain paired child states and select state pairs with value differences larger than a specified threshold. After obtaining preference data, we perform process-based preference learning using direct preference optimization (DPO) Rafailov et al. (2024) to self-improve the TQA model.
@@
Experiments show that ppt effectively improves TQA models by up to 5% on in-domain datasets and 2.4% on out-of-domain datasets with only 8k preference data. It achieves similar performances compared to methods requiring additional LLMs as judges, yet being ten times more efficient. The enhanced model outperforms existing fine-tuned TQA models. On three datasets, it even achieves similar performances compared with much larger and complex frameworks, yet boasts five times more efficiency during inference.
2 Related Work
We discuss current approaches in addressing TQA and methods for self-improvements.
Table Question Answering.
Current TQA systems can be categorized into two groups: fine-tuned models Zhang et al. (2023a); Wu and Feng (2024) and training-free frameworks Zhou et al. (2025); Nahid and Rafiei (2024). The former collects existing TQA datasets to train an end-to-end system that either directly outputs answers Zhang et al. (2023a) or makes use of an additional reasoning chain Wu and Feng (2024). small-size models ($\leq 8B$) are utilized out of efficiency considerations. Training-free frameworks in TQA are dependent on larger models, e.g., GPT-3.5, Qwen-72B Qwen et al. (2025) and involve more complex designs, such as tools Nahid and Rafiei (2024) and agentic collaborations Zhou et al. (2025). We focus on the fine-tuning approach given its inference efficiency.
Self-Improving LLMs.
Self-improving focuses on using training data generated by the model itself to further obtain performance gains Singh et al. (2024); Zelikman et al. (2024); Xiong et al. (2024); He et al. (2024). Common methods applied for fine-tuning include rejected sampling fine-tuning (RST) Yuan et al. (2023) and DPO Tu et al. (2025); Xu et al. (2024); Zhang et al. (2025a). To obtain step-DPO data for preference learning, previous work either uses closed-source models for locating errors Lai et al. (2024) or existing process supervision datasets Lightman et al. (2023) to train a step reward model. Zhang et al. (2025a). In contrast, our framework does not involve LLM judges, nor does it depend on manually annotated process supervision data. To the best of our knowledge, we are the first to explore process-based supervision for self-improvement in TQA.
3 Process-based Preference Learning for TQA
Task Setup.
Given a table $t$, a question $q$, a fine-tuned TQA model $M_{ft}$ should provide a correct answer $a$. To enhance systems’ performances, a reasoning chain $r$ is generated along with $a$. $r$ consists of $m$ steps: $\{k_{1},k_{2},...,k_{m}\}$. A problem ($t,q$) and an instruction ($u$) form an initial state $s_{0}$. Each new state $s_{i}$ consists of previous state $s_{i-1}$ and a step $k_{i}$ generated at the timestep $i$: $s_{i}=(s_{i-1},k_{i})$ and $s_{i-1}=\{s_{0},k_{1},k_{2},....k_{i-1}\}$.
State Generation.
Given a TQA problem, we first sample $m$ reasoning chains from $M_{ft}$, resulting in $\{r\}_{i=1}^{m}$. Prompts can be found in Appendix A.1. Each reasoning chain consists of multiple steps. We accumulate steps to construct states: $s_{i}=\{s_{0},k_{1},...k_{i}\}$. If all reasoning traces lead to correct answers, we abandon the problem as it might be too easy and only lead to high value states.
@@ 3 Process-based Preference Learning for TQA
A state value function $V$ takes in a state and returns its value. We approximate a state’s value by Monte Carlo (MC) sampling: $V(s_{i})$ is calculated as the probability of $s_{i}$ leading to the correct answer. More specifically, $M_{ft}$ takes in $s_{i}$ and completes the current reasoning chain until reaching an answer. This is repeated $n$ times, resulting in $\{(s_{i+1,j},...,s_{z_{j},j},a_{j})\}_{j=1}^{n}$, where $a_{j}$ and $z_{j}$ are the predicted answer and the total number of steps for the $j$-th finalized solution, respectively. $V(s_{i})$ can be represented as:
$V(s_{i})=\frac{\sum_{j=1}^{n}\mathbb{I}(a_{j}=a)}{n}$ (1)
The continuous value allows more flexible selection of pair-wise steps described in the following paragraph as we can control the distance between the selected states.
Pair-wise Step Sampling
After obtaining values for each state, we filter out intermediate states $s_{i}$ where $V(s_{i})=0$ and also their child state $\{s_{i+1},...s_{z}\}$. This is exemplified by the red nodes in Figure 2. We assume a state of value 0 to contain erroneous steps. Rolling out from it is likely to create bad quality child states. For each remaining state $s_{i}$, we use the completion traces sampled when calculating $V(s_{i-1})$: $\{(s_{i,j},...,s_{z_{j},j},a_{j})\}_{j=1}^{n}$ as roll outs, and calculate state values for each sampled $\{s_{i,j}\}_{j=1}^{n}$. This results in a set of pair-wise states: $(s_{i},\{s_{i,j}\}_{j=1}^{n})$ that can be used to construct step-wise preference dataset $D_{sdpo}$. As $V(s_{i})$ is a continuous value, a good state $s_{w}$ and a bad state $s_{l}$ pair is selected if $V(s_{w})-V(s_{l})\geq\tau$, where $\tau$ is a hyper-parameter. The preference data for step DPO can be represented as $D_{sdpo}=\{(s_{i-1},k_{i}^{w},k_{i}^{l})^{d}\}_{d=1}^{|D_{sdpo}|}$.
Preference Training.
After gathering the process-based preference dataset, we fine-tune $M_{ft}$ using pairs of good and bad steps given previous steps. The loss function is:
$\mathcal{L}_{\text{SDPO}}=-\mathbb{E}_{(s_{i-1},k_{i}^{w},k_{i}^{l})\sim\mathcal{D}_{sdpo}}\left[\log\sigma\left(\beta\log\frac{\pi_{\theta}(k_{i}^{w}|s_{i-1})}{\pi_{ref}(k_{i}^{w}|s_{i-1})}\right)\right.\\ \left.-\beta\log\frac{\pi_{\theta}(k_{i}^{l}|s_{i-1})}{\pi_{ref}(k_{i}^{l}|s_{i-1})}\right]\end{aligned}$
@@ 4 Experiments
Note that existing fine-tuned TQA systems do not feature clear step separation in their reasoning chains. Separating steps by line breaks as in math Luo et al. (2024) is problematic since a line break also indicates a table row in TQA. Therefore, we fine-tune TQA models ourselves that can output clearly separated steps. Details are presented in Appendix A.3. We term the fine-tuned model $M_{ft}$.
@@
We consider baselines for self-improvement strategies: RFT Yuan et al. (2023) trains a model with correct reasoning traces using supervised fine-tuning. FDPO Xu et al. (2024) trains a model with pair-wise correct and incorrect full reasoning chains using DPO. Baselines for value functions include: MC with binary labels (mc-b), it returns binary state values based on if states derive final correct answers. Mixed estimation (mix) Zhang et al. (2025c) scores $s_{i}$ 1 if both mc-b and an external LLM judge $M_{j}$ output 1. If both judges return 0, $s_{i}$ is 0. States receiving different scores by the judges are not considered for building the preference dataset. This accounts for approximately 36% of data (Figure 5). Baselines for TQA models include both end-to-end Zhang et al. (2023a); Wu and Feng (2024) and training-free frameworks Zhou et al. (2025); Nahid and Rafiei (2024). Details of the base methods and models can be found in Appendix A.4.
Figure 2: Process-based preference data collection. We estimate a state value by the probability of a state leading to a correct answer. In the first example, $V(s_{i})=\frac{2}{3}$. After obtaining state values, we do not consider intermediate states that have a value of 0 ($s_{21}$), together with their child states ($s_{31}$). We sample pair-wise states for each remaining state, e.g., $s^{\prime}_{22}$ is sampled by rolling out $s_{12}$ and is regarded as a pair state for $s_{22}$.
@@
We train $M_{ft}$ using the training sets of three common TQA datasets: WTQ Pasupat and Liang (2015), TabFact Chen et al. (2020), and HiTab Cheng et al. (2022). To obtain preference data, we sample from their validation sets. We use the test sets of these three datasets as in-domain evaluation data, and incorporate three out-of-domain datasets: WikiSQL (Zhong et al., 2017), SCITAB Lu et al. (2023), and CRT Zhang et al. (2023b) to test models’ generalisability. Details about the datasets are presented in Appendix A.2.
@@
We choose Qwen-2.5-7B Qwen et al. (2025) and LlaMA-3.1-8B Grattafiori et al. (2024) as base models. During preference learning, we fix the fine-tuning dataset size to 8k for Qwen-2.5-7B and 6.7k for LlaMA-3.1-8B, as different baselines result in different sample sizes (statistics are shown in Appendix A.5). For fair comparison, we use the smallest sample size collected as the fine-tuning data size. We make sure all sampled data correspond to the same set of TQA problems to rule out instance-level effects on fine-tuning. Hyper-parameters are shown in Appendix A.6. We use Qwen-2.5-72B as $M_{j}$. The Prompt for $M_{j}$ is shown in Appendix A.1. The number of reasoning chains $m$ is set to 4 and the roll-out number $n$ is set to 8. The threshold $\tau$ is set to 0.9 and temperature is set to 0.7 and 0 during dataset construction and inference respectively. We use Exact Match as the evaluation metric. All experiments are conducted using 4 A100 GPUs. Training is performed with LlaMA-Factory Zheng et al. (2024) and inference is performed with VLLM Kwon et al. (2023).

### v2

# p2-TQA: A Process-based Preference Learning Framework for Self-Improving Table Question Answering Models

@@ Abstract
Table question answering (TQA) focuses on answering questions based on tabular data. Developing TQA systems targets effective interaction with tabular data for tasks such as cell retrieval and data analysis. While recent work has leveraged fine-tuning to improve TQA systems, existing approaches often under-utilize available data and neglect the potential of post-training for further gains. In this work, we introduce p2-TQA, a process-based preference learning framework for TQA post-training. p2-TQA automatically constructs process-based preference data via a table-specific pipeline, eliminating the need for manual or costly data collection. It then optimizes models through contrastive learning on the collected data. Experiments show that p2-TQA effectively improves TQA models by up to 5% on in-domain datasets and 2.4% on out-of-domain datasets with only 8,000 training instances. Furthermore, models enhanced with p2-TQA achieve competitive results against larger, more complex state-of-the-art TQA systems, while maintaining up to five times higher efficiency.
@@ 1 Introduction
Table question answering (TQA) aims to generate accurate responses to queries over tables. Current TQA systems fall into two categories: fine-tuned models zhang2025tablellmenablingtabulardata; wu-feng-2024-protrix and training-free frameworks zhou-etal-2025-efficient; nahid-rafiei-2024-tabsqlify. The former fine-tune pre-trained small-size ($\leq$ 8B) large language models (LLMs) while the latter rely on large LLMs and involve complex designs. There is a growing interest in understanding and developing fine-tuned TQA models deng2025betterunderstandingtableinstruction; deng2025rethinkingtableinstructiontuning due to their promising performance and inference efficiency.
Current fine-tuning methods for TQA often augment only a subset of existing datasets with Chain-of-Thought (CoT) reasoning wei2023chainofthoughtpromptingelicitsreasoning, constrained by the high cost of querying large commercial models wu-feng-2024-protrix. Models are then trained on these reasoning chains and answers via supervised fine-tuning. This workflow has two key limitations: (1) only part of the training data is used, i.e., data is under-utilized; and (2) potential performance gains from post-training are neglected. Based on the identified gaps, this study aims to answer the research question: How can we leverage existing datasets to post-train a TQA model for further performance gains? Our effective post-training method (shown in Figure 1) enhances models without requiring additional manual or costly data.
Figure 1: An overview of p2-TQA: An existing model generates reasoning chains for a given problem. The chains are parsed into states, composed of cumulative steps. Each state is scored by a value function. We then create pairwise steps by rolling out parent states, selecting those with value differences exceeding a threshold. Lastly, contrastive learning is performed over collected data to improve the TQA model.
While post-training with self-generated data has primarily been explored in mathematics and coding singh2024humandatascalingselftraining; zelikman2024selftaughtoptimizerstoprecursively; xiong-etal-2024-watch; he2024semisupervisedrewardmodelingiterative, particularly through step-wise preference learning tu2025enhancingllmreasoningiterative; xu-etal-2024-chatglm, its application in TQA remains under-explored. Previous methods for obtaining step-wise preference pairs typically rely on (1) closed-source or large open-source models as judges to discern correct and incorrect steps lai2024stepdpostepwisepreferenceoptimization, (2) Monte Carlo Sampling (MCS) to estimate a step’s quality wang-etal-2024-math; xiong-etal-2024-watch; hwang-etal-2024-self, or (3) a combination of both zhang2025lessonsdevelopingprocessreward. However, performing step-wise preference learning to TQA presents unique challenges. Compared to mathematics, it typically involves longer inputs and intermediate reasoning due to large tables. This necessitates a careful design of steps; simply using newline breaks to obtain steps, as in math, could lead to excessive computation costs, because newline breaks also denote new rows in tables. Furthermore, the structured nature of input in TQA demands a reconsideration of value functions, since LLM judges have a limited understanding of table structures sui2024tablemeetsllmlarge, especially with larger tables zhou-etal-2024-freb.
To this end, we introduce p2-TQA, a process-based preference learning pipeline for TQA post-training: p2-TQA operates in three stages for data collection (Figure 1): state generation, state value estimation, and pair-wise step sampling. The first stage collects and parses reasoning chains into states, composed of cumulative reasoning steps that are carefully designed for TQA. The last two stages construct step-wise preference pairs via MCS for state value estimation and a stringent filtering process for quality control. We apply direct preference optimization (DPO) rafailov2024directpreferenceoptimizationlanguage with the collected data to self-improve a TQA model.
@@
Experiments show that p2-TQA improves TQA models by up to 5% on in-domain datasets and by up to 2.4% on out-of-domain datasets with only 8k preference training pairs. It surpasses methods that require additional LLMs as judges, yet it is ten times more efficient. This underscores our contribution in establishing an effective and efficient framework for self-improving TQA models. The self-improved models outperform existing fine-tuned TQA models and achieve comparable performance to much larger and more complex frameworks on three datasets, while maintaining five times higher inference efficiency. Code is available.  [^1] https://github.com/boschresearch/p2-TQA
2 Step-wise Preference Learning for TQA
Given a table $t$, a question $q$, and a fine-tuned TQA model $M_{ft}$ that outputs a reasoning chain $r$ consisting of $l$ steps: $\{k_{1},k_{2},...,k_{l}\}$, along with a predicted answer $a$, our goal is to collect high-quality step-wise preference data $(k_{i}^{good},k_{i}^{bad})$ from $M_{ft}$ and perform contrastive learning to improve $M_{ft}$.
Step Design and State Generation.
Designing an effective step scope is crucial for both performance and sampling efficiency. Inspired by SQL query operations, we define a step as a basic operation, such as filtering or counting. To aid LLM reasoning in TQA, each step includes a planning component that outlines the required operation and a reasoning component that provides its result, e.g., Count the number of gold medals received in 2004. There are 6 gold medals received in 2004. For each TQA problem, an initial state $s_{0}$ is formed from $t$, $q$, and an instruction $u$. We then sample $m$ reasoning chains, denoted as $\{r\}_{i=1}^{m}$ from $M_{ft}$. Prompts can be found in Appendix A.1. A new state $s_{i}$ is defined as the combination of previous state $s_{i-1}$ and a step $k_{i}$ generated at the timestep $i$: $s_{i}=(s_{i-1},k_{i})$, where $s_{i-1}=\{s_{0},k_{1},k_{2},....k_{i-1}\}$. Problems where all reasoning traces lead to correct answers are discarded, as they are considered too easy.
@@ 2 Step-wise Preference Learning for TQA
A state value function $V$ takes in a state and returns its value. We approximate a state’s value by using Monte Carlo Sampling similar to wang-etal-2024-math: $M_{ft}$ takes in $s_{i}$ and completes the current reasoning chain until reaching an answer. This is repeated $n$ times. $V(s_{i})$ is calculated as the probability of $s_{i}$ leading to the correct answer. An example is shown in Figure 2, where $V(s_{i})=\frac{2}{3}$. The continuous value allows a flexible and controlled selection of pair-wise steps described in the following paragraph.
Figure 2: Process-based preference data collection. We estimate a state value by the probability of a state leading to a correct answer. In the first example, $V(s_{i})=\frac{2}{3}$. After obtaining state values, we do not consider intermediate states that have a value of 0 ($s_{21}$), together with their child states. We sample pair-wise states for each remaining state, e.g., $s^{\prime}_{22}$ is sampled by rolling out $s_{12}$ and is regarded as a pair state for $s_{22}$.
Pair-wise Step Sampling.
After obtaining state values, we filter out intermediate states $s_{i}$, where $V(s_{i})=0$, and also remove their child states $\{s_{i+1},...s_{z}\}$. This is exemplified by the red nodes in Figure 2. We assume a state of value 0 to contain erroneous steps. Rolling out from it is likely to create bad-quality child states. For each remaining state $s_{i}$, we use the completion traces sampled when calculating $V(s_{i-1})$: $\{(s_{i,j},...,s_{z_{j},j},a_{j})\}_{j=1}^{n}$ as rollouts, where $a_{j}$ is the predicted answer and $z_{j}$ is the total number of steps for the $j$-th finalized solution. Next, we calculate state values for each sampled $\{s_{i,j}\}_{j=1}^{n}$. This results in a set of pair-wise states: $(s_{i},\{s_{i,j}\}_{j=1}^{n})$ that can be used to construct step-wise preference dataset $D_{sdpo}$. As $V(s_{i})$ is a continuous value, a pair comprising one good state $s_{good}$ and one bad state $s_{bad}$ is selected if $V(s_{good})-V(s_{bad})\geq\tau$, where $\tau$ is a hyper-parameter. We prove later in our experiments that for TQA, this filtering mechanism greatly improves performance and efficiency on top of using MCS as the value function, while maintaining efficiency. The preference data for step DPO can be represented as $D_{sdpo}=\{(s_{i-1},k_{i}^{good},k_{i}^{bad})^{d}\}_{d=1}^{|D_{sdpo}|}$.
After collecting the preference dataset, we fine-tune $M_{ft}$ using pairs of good and bad steps given previous steps. The loss function is defined as follows, where $\beta$ is a hyper-parameter controlling the strength of incorporating the preference signal. $\pi_{ref}$ and $\pi_{\theta}$ denote the original reference and updated model, respectively.
$\begin{aligned} \mathcal{L}=-\mathbb{E}_{(s_{i-1},k_{i}^{good},k_{i}^{bad})\sim\mathcal{D}_{sdpo}}\left[\log\sigma\left(\beta\log\frac{\pi_{\theta}(k_{i}^{good}|s_{i-1})}{\pi_{ref}(k_{i}^{good}|s_{i-1})}\right)\right.\\ \left.-\beta\log\frac{\pi_{\theta}(k_{i}^{bad}|s_{i-1})}{\pi_{ref}(k_{i}^{bad}|s_{i-1})}\right]\end{aligned}$
@@ 3 Experiments
We present details for TQA models, baselines, datasets, and experimental settings in this section.
@@
Existing fine-tuned TQA models do not feature clear step separations. We therefore obtain a TQA model $M_{ft}$ by fine-tuning an LLM using the step definition introduced before. Following previous work wu-feng-2024-protrix; zhang2025tablellmenablingtabulardata, we employ Deepseek-V3 deepseekai2025deepseekv3technicalreport to generate reasoning chains. We sample 2.4k, 1.5k, and 2.3k examples from the training sets of WTQ pasupat-liang-2015-compositional, TabFact 2019TabFactA, and HiTab cheng-etal-2022-hitab, respectively. We prompt Deepseek-V3 to produce reasoning chains along with final answers, retaining only those chains that yield correct answers. This process results in 1,612 instances from WTQ, 1,425 from TabFact, and 1,277 from HiTab, for a total of 4,314 instances.
@@
We consider the following baselines for self-improvement strategies: (1) RFT yuan2023scalingrelationshiplearningmathematical trains a model with self-generated reasoning traces that lead to correct answers using supervised fine-tuning. (2) FDPO xu-etal-2024-chatglm trains a model with pair-wise correct and incorrect full reasoning chains using DPO.
Baselines for value functions include: (3) MC with binary labels (mc-b) wang-etal-2024-math returns binary state values based on whether states derive final correct answers. Mixed estimation (mix) zhang2025lessonsdevelopingprocessreward scores $s_{i}$ 1 if both mc-b and an external LLM judge $M_{j}$ output 1. If both judges return 0, $s_{i}$ is 0. States receiving different scores from the judges are not considered for building the preference dataset. (4) Self-Explore hwang-etal-2024-self randomly selects a preferred reasoning trace and uses full completion of it instead of a step. This results in longer preferred responses over rejected ones.
Baselines for TQA models include both end-to-end and training-free frameworks: (1) TableLlaMA Zhang2023TableLlamaTO is an end-to-end fine-tuned model with LlaMA-2-7B Touvron2023Llama2O as the base model. (2) Protrix wu-feng-2024-protrix is fine-tuned with around 4k instances with reasoning chains generated from GPT-4, also using LlaMA-2-7B as the base model. (3) MACT zhou-etal-2025-efficient is a training-free framework, leveraging tools and agent collaboration. (4) TabSQLify nahid-rafiei-2024-tabsqlify decomposes tables into relevant sub-tables with SQL query generation and execution. Then, sub-tables and questions are passed to LLMs to obtain final answers.
@@
We train $M_{ft}$ using the training sets of WTQ, TabFact, and HiTab. To obtain preference data, we sample from their validation sets. We use the test sets of these three datasets as in-domain evaluation data and incorporate three out-of-domain datasets: WikiSQL (zhongSeq2SQL2017), SCITAB lu-etal-2023-scitab, and CRT zhang-etal-2023-crt to test models’ generalisability. These datasets test our method in various degrees of complexity. Thus, we make sure our method is generally effective. Details about the datasets are presented in Appendix A.2.
@@
We choose Qwen-2.5-7B qwen2025qwen25technicalreport and LlaMA-3.1-8B grattafiori2024llama3herdmodels as base models. During preference learning, we fix the fine-tuning dataset size to 8k for Qwen-2.5-7B and 6.7k for LlaMA-3.1-8B, as different baselines result in different sample sizes (statistics are shown in Appendix A.3). For fair comparison, we use the smallest sample size collected as the fine-tuning data size. Hyper-parameters are shown in Appendix A.4. We use Qwen-2.5-72B as $M_{j}$ (Prompt in Appendix A.1). The number of reasoning chains $m$ is set to 4, and the roll-out number $n$ is set to 8. The threshold $\tau$ is set to 0.9, and the temperature is set to 0.7 and 0 during dataset construction and inference, respectively. We use Exact Match (EM) as the evaluation metric. All experiments are conducted using 4 A100 GPUs. Training is performed with LlaMA-Factory zheng2024llamafactory and inference is performed with VLLM kwon2023efficient.